The gas mileage G(s) (in miles/gallon) of a Ford Tauras when driven at a speed of S mph can be approximated by the function
G(s)=-0.02s^2+2s-25, 25 - -
a. Find the gas mileafe if the car is driven at a speed of 60 mph.
b. At what speed(s) does the car get 17 mpg.

Answer by ankor@dixie-net.com(22740)

The gas mileage G(s) (in miles/gallon) of a Ford Tauras when driven at a speed of S mph can be approximated by the function
G(s)=-0.02s^2+2s-25, 25 - -
a. Find the gas mileage if the car is driven at a speed of 60 mph.
G(s) = -.02s^2 + 2s - 25
when s = 60
G(s) = -.02(60^2) + 2(60) - 25
G(s) = -.02(3600) + 120 - 25
G(s) = -72 + 120 - 25
G(s) = 23 mpg at 60 mph
:
b. At what speed(s) does the car get 17 mpg.
G(s) = 17
-.02s^2 + 2s - 25 = 17
-.02s^2 + 2s - 25 - 17 = 0
-.02s^2 + 2s - 42 = 0
Use the quadratic formula; a=-.02; b=2; c=-42
I got two solutions
s = 30 mph
and
s = 70 mph, is the solution that makes the most sense, since going 10 mph faster would probably lower the mpg from 23 to 17